# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ 3.2 Conformal inference via locally exchangeable subgroups
where
$Z_{\ell}^{{\textnormal{swap}(i,j)}}(y)=\begin{cases}Z_{n+j}(y)&\ell=i,\\ Z_{\ell}&\ell\neq i.\end{cases}\quad X_{n+\ell}^{{\textnormal{swap}(i,j)}}=\begin{cases}X_{i}&\ell=j,\\ X_{n+\ell}&\ell\neq j.\end{cases}$
@@
That is, they are the calibration and test data if we treat $Z_{i}$ as the $j$-th test point, and $Z_{n+j}(y)$ as the $i$-th calibration point. Figure 1 is an illustration of the swapping operation. Applying the same selection rule, the selection set after swapping and with the hypothesized $y$ is then
@@
^$\widehat{{\mathcal{R}}}_{n+j}(y)=\big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}\colon j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(y),\text{ and }\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(y)\in\mathfrak{S}\big{\}}.$ (7)
In words, the reference set is the collection of calibration points $i$ such that, after swapping unit $i$ and $n+j$, the test point $j$ remains in the focal set and the focal set remains in $\mathfrak{S}$. We use the notation ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ to emphasize that ^$\widehat{\mathcal{R}}_{n+j}(\cdot)$ is a data-dependent mapping from ${\mathcal{Y}}$ to subsets of ${\mathcal{I}}_{{\textnormal{calib}}}$.
With all the preparation, we define our prediction set via
@@
As we shall show shortly, our prediction set (8) achieves near-exact coverage when the nonconformity score is continuous and the reference set is of moderate size. In fact, it is possible to achieve exact coverage by introducing extra randomness:
@@
=$\quad=\bigg{\{}y\colon\frac{\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)<V_{i}\}+U_{j}\cdot(1+\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)=V_{i}\})}{1+|\widehat{\mathcal{R}}_{n+j}(y)|}\leq 1-\alpha\bigg{\}},$ (9)
where $U_{1},\ldots,U_{m}$ are i.i.d. random variables drawn from $\textrm{Unif}[0,1]$ which are independent of the data.
Remark 1 (Connection to MCP).
Our framework is closely related to Mondrian Conformal Prediction (MCP(Vovk et al.,, 2003) which applies to one test unit ($m=1$). Given $K$ categories $[K]$, they assume access to a classifier ${\mathcal{K}}$ that gives the class label for all units: $(\kappa_{1},\kappa_{2},\dots,\kappa_{n+1})={\mathcal{K}}(Z_{1},Z_{2},\dots,Z_{n+1}),$ where $\kappa_{i}\in[K]$ is the class label for unit $i$. The classifier ${\mathcal{K}}$ is required to be invariant to any permutation $\pi$ of $[n+1]$, i.e., $(\kappa_{\pi(1)},\kappa_{\pi(2)},\dots,\kappa_{\pi(n+1)})={\mathcal{K}}(Z_{\pi(1)},Z_{\pi(2)},\dots,Z_{\pi(n+1)}).$ The goal of MCP is to construct a prediction set for the test unit with cateogory-wise coverage, i.e., $\mathbb{P}(Y_{n+1}\in\widehat{C}^{\textnormal{MCP}}_{\alpha,n+1}{\,|\,}\kappa_{n+1}=k)$ for each $k\in[K]$. This is achieved by finding a subset of calibration data for each hypothesized $y\in{\mathcal{Y}}$,
~$\widetilde{{\mathcal{R}}}(y)=\{i\in{\mathcal{I}}_{\textnormal{calib}}:{\mathcal{K}}({\mathcal{D}}_{{\textnormal{calib}}},(X_{n+1},y))_{[i]}=k\},$ (10)
and then constructing the prediction set as
While the setups seem different, JOMI effectively reduces to MCP when ${\mathcal{K}}$ is a binary classifier and ${\mathcal{S}}$ is equivariant under permutations of $[n+1]$. To see this, one can view the selection rule ${\mathcal{S}}$ as categorizing test points into two classes (selected and not selected), and selection-conditional coverage (3) as coverage given the “selected” category. In this way, one can check that our reference set ^$\widehat{{\mathcal{R}}}_{n+1}(y)$ coincides with ~$\widetilde{{\mathcal{R}}}(y)$ defined in (10), and thus the prediction sets also coincide.
With a distinct goal of selection-conditional coverage, we additionally address the challenge that realistic selection rules are often asymmetric to calibration and (multiple) test points. To this end, JOMI generalizes MCP via a more delicate but still general framework for finding exchangeable subgroups. As we jointly consider the calibration and test points when defining the reference set, we call our method JOint Mondrain Conformal Inference (JOMI).
3.3 Theoretical guarantee
Theorem 1 confirms the validity of our prediction sets (8) and (3.2).
@@ 3.3 Theoretical guarantee
Suppose ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is invariant to permutations of ${\mathcal{D}}_{\textnormal{calib}}$. Suppose $\{Z_{i}\}_{i=1}^{n}\cup\{Z_{n+j}\}$ are exchangeable conditional on $\{X_{n+\ell}\}_{\ell\in[m]\backslash\{j\}}$ for any $j\in[m]$. Then, for any selection taxonomy $\mathfrak{S}$,
@@
    Furthermore, if ties between $V_{1},\ldots,V_{n},V_{n+j}$ occur with probability zero, then
@@
We defer the detailed proof of Theorem 1 to Appendix B.4. Let us provide some intuitions for the validity results here. Similar to the ideas of SCP in Section 3.1, the key fact we rely on is that, plugging in the true value $y=Y_{n+j}$, data in the reference set and the new test point are still exchangeable conditional on the event that $j$ is selected. To be specific, imagine that we know the unordered set $[Z_{1},\dots,Z_{n},Z_{n+j}]=[z_{1},\dots,z_{n},z_{n+j}]$ and the values of the remaining test points $\{X_{n+\ell}\}_{\ell\neq j}$. The only randomness is thus in the ordering of $Z_{1},\dots,Z_{n+j}$. Meanwhile, our reference set ^$\widehat{\mathcal{R}}_{j}(\cdot)$ is constructed in a delicate way such that the (unordered) scores in the reference set $[V_{i}\colon i\in\widehat{\mathcal{R}}_{j}(Y_{n+j})]$ is fully determined by the unordered set $[Z_{1},\dots,Z_{n},Z_{n+j}]$ if ${\mathcal{S}}(\cdot,\cdot)$ is invariant to permutations of the “calibration” data in the first argument. Then, by the exchangeability of $Z_{1},\dots,Z_{n},Z_{n+j}$, the probability of $Y_{n+j}$ taking on any value in $\widehat{R}$ is equal given the unordered set, leading to the validity in Theorem 1.
@@ 4 Computationally tractable instances
Section 3 provides a general framework for constructing valid prediction sets conditional on selection events, but computing the prediction sets according to their definitions requires looping over all possible values of $y$, often a computationally intractable task.
In this section, we instantiate the generic procedure through concrete examples, in which special structures allow us to efficiently compute the prediction sets (8) and (3.2). We focus on three classes of selection rules: the selection rules using solely the covariates, those based on conformal p-values, and those based on conformal prediction sets.
@@ 4.1 Covariate-dependent selection rules
When the selection rule depends solely on the covariates, the reference sets can be efficiently computed by definition, since ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ no longer depends on $y$ (we shall also suppress the dependence of ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ and $S(x,y)$ on $y$ throughout Section 4.1). The complete procedure for constructing $\widehat{C}_{\alpha,n+j}$’s and $\widehat{C}^{\textnormal{rand}}_{\alpha,n+j}$ with an arbitrary covariate-dependent selection rule is summarized in Algorithm 1, and the computational complexity the procedure is $O(mn\cdot C_{\mathcal{S}})$, with $C_{\mathcal{S}}$ being the complexity of the selection process.
Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$;
selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$;
@@
for $j=1,\ldots,m$ do
       /* Constructing reference sets */
       Initialize ^$\widehat{{\mathcal{R}}}_{n+j}=\emptyset$;
       for $i=1,\ldots,n$ do
             ^$\widehat{{\mathcal{R}}}_{n+j}\leftarrow\widehat{{\mathcal{R}}}_{n+j}\cup\{i\}$   if $j\in\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}$ and $\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}$;
       end for
       /* Constructing prediction sets */
       if form = dtm then
             $\widehat{C}_{\alpha,n+j}=\Big{\{}y\in{\mathcal{Y}}\colon V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}}\cup\{+\infty\}\big{)}\Big{\}}.$
       end if
      if form = rand then
             Sample $U_{j}\sim\textnormal{Unif}[0,1]$;
             $\widehat{C}_{\alpha,n+j}=\bigg{\{}y\colon\frac{\sum_{i\in\widehat{\mathcal{R}}_{n+j}}\mathds{1}\{V(X_{n+j},y)<V_{i}\}+U_{j}\cdot(1+\sum_{i\in\widehat{\mathcal{R}}_{n+j}}\mathds{1}\{V(X_{n+j},y)=V_{i}\})}{1+|\widehat{\mathcal{R}}_{n+j}|}\leq 1-\alpha\bigg{\}}.$
       end if
end for
Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
Algorithm 1 JOMI for arbitrary covariate-dependent selection rules
By Theorem 1, the output of Algorithm 1 is valid as long as the selection rule is invariant to the permutation of the calibration points. This includes many commonly-used selection rules:
@@
    Top-K selection: the $K$ test units with the highest scores $S(X_{i})$ are selected.
@@
    Selection based on joint quantiles: unit $i$ is selected if its score $S(X_{i})$ is surpasses the $q$-th quantile of both the calibration and test scores.
@@
    Selection based on calibration quantiles: a test unit $i$ is selected if its score $S(X_{i})$ surpasses the $q$-th quantile of calibration scores.

### v3

@@ 3.2 Conformal inference via a reference set
where for $k\in[n]$ and $\ell\in[m]$,
$Z_{k}^{{\textnormal{swap}(i,j)}}(y)=\begin{cases}Z_{n+j}(y)&k=i,\\ Z_{k}&k\neq i.\end{cases}\quad X_{n+\ell}^{{\textnormal{swap}(i,j)}}=\begin{cases}X_{i}&\ell=j,\\ X_{n+\ell}&\ell\neq j.\end{cases}$
That is, ${\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}(y)$ and ${\mathcal{D}}_{\textnormal{test}}^{\textnormal{swap}(i,j)}$ are the calibration and test data if we treat $Z_{i}$ as the $j$-th test point, and $Z_{n+j}(y)$ as the $i$-th calibration point. Figure 2 is an illustration of the swap operation.
@@
Applying the same selection rule ${\mathcal{S}}$ to the swapped data, we define the swapped selection set with the hypothesized $y$ as
@@
^$\widehat{{\mathcal{R}}}_{n+j}(y)=\big{\{}i\in[n]\colon j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(y),\text{ and }\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(y)\in\mathfrak{S}\big{\}}.$ (8)
In words, the reference set is the collection of calibration points $i\in[n]$ such that, after swapping unit $i$ and $n+j$, the (posited) test point $j$ (which is the original unit $i$) remains in the focal set and the focal set remains in $\mathfrak{S}$. We use the notation ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ to emphasize that ^$\widehat{\mathcal{R}}_{n+j}(\cdot)$ is a data-dependent mapping from ${\mathcal{Y}}$ to the power set of $[n]$.
With all the preparation, we define our prediction set for $Y_{n+j}$ as
@@
As we shall show shortly, our prediction set (9) achieves near-exact coverage when the nonconformity score is continuous and the reference set is of moderate size. In addition, we can achieve exact coverage by introducing extra randomness:
@@
$=$ $\bigg{\{}y\colon\frac{\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)<V_{i}\}+U_{j}\cdot(1+\sum_{i\in\widehat{\mathcal{R}}_{n+j}(y)}\mathds{1}\{V(X_{n+j},y)=V_{i}\})}{1+|\widehat{\mathcal{R}}_{n+j}(y)|}\leq 1-\alpha\bigg{\}},$ (10)
where $U_{1},\ldots,U_{m}$ are i.i.d. random variables drawn from $\textrm{Unif}[0,1]$ independent of the data. In Supplementary Sections A.2 and A.3, we discuss in detail the connection of our method to Mondrian conformal prediction (MCP) and comparison with the BY procedure (Benjamini and Yekutieli,, 2005). While the BY procedure is a natural and heuristic solution to post-selection inference, it suffers from over-conservativeness and non-adaptivity to the selection event.
3.3 Theoretical guarantees
Theorem 1 confirms the conditional validity of our prediction sets $\widehat{C}_{\alpha,n+j}$ in (9) and $\widehat{C}^{\textnormal{rand}}_{\alpha,n+j}$ in (3.2).
@@ 3.3 Theoretical guarantees
Suppose ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is invariant to permutations of ${\mathcal{D}}_{\textnormal{calib}}$, and that $\{Z_{i}\}_{i=1}^{n}\cup\{Z_{n+j}\}$ are exchangeable conditional on $\{X_{n+\ell}\}_{\ell\in[m]\backslash\{j\}}$ for any $j\in[m]$. Then, for any selection taxonomy $\mathfrak{S}$, the following statements hold.
@@
    Furthermore, if ties among $V_{1},\ldots,V_{n},V_{n+j}$ occur with probability zero, then
@@
We defer the detailed proof of Theorem 1 to Supplementary Section D.4 and provide some intuition here. Similar to the ideas of SCP in Section 3.1, the key fact we rely on is that, plugging in the true value $y=Y_{n+j}$, data in the reference set and the new test point are still exchangeable conditional on the selection event. To be specific, to prove (11), we are to show a stronger result:
$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{V_{n+j}\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S},\,[{\mathcal{D}}_{j}],\,{\mathcal{D}}_{j}^{c}\Big{)}\geq 1-\alpha,$ (13)
where we recall $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$ and ${\mathcal{D}}_{j}^{c}={\mathcal{D}}_{\textnormal{test}}\backslash\{X_{n+j}\}$. For any fixed values $z_{1},\dots,z_{n},z_{n+j}$, once given the unordered values $[{\mathcal{D}}_{j}]=[d_{j}]=[z_{1},\dots,z_{n},z_{n+j}]$ and the values of other test points $\{X_{n+\ell}\}_{\ell\neq j}$, the only randomness is in the ordering of $Z_{1},\dots,Z_{n+j}$ among $[z_{1},\dots,z_{n},z_{n+j}]$. Meanwhile, we show that our reference set ^$\widehat{\mathcal{R}}_{n+j}(\cdot)$ is constructed in a delicate way such that the (unordered set of) scores $\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{V_{n+j}\}$ is fully determined by $[z_{1},\dots,z_{n},z_{n+j}]$. That is, $\widehat{R}_{n+j}^{+}:=[V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}]$ is fully determined given $[d_{j}]$. Then, (13) reduces to
$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\widehat{R}_{n+j}^{+}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S},\,[{\mathcal{D}}_{j}],\,{\mathcal{D}}_{j}^{c}\Big{)}\geq 1-\alpha,$ (14)
Finally, by the exchangeability of $Z_{1},\dots,Z_{n},Z_{n+j}$, the probability of $V_{n+j}$ taking on any value in $\widehat{R}_{n+j}^{+}$ is equal given $[{\mathcal{D}}_{j}]$ and ${\mathcal{D}}_{j}^{c}$, leading to the validity in Theorem 1 via Bayes’ rule.
@@ 4 Computationally tractable instances
So far, we have presented a general framework for constructing valid prediction sets conditional on general selection events. However, computing the prediction sets according to their definition requires looping over all possible values of $y\in{\mathcal{Y}}$, which can be computationally intractable. When $|{\mathcal{Y}}|$ is finite (and relatively small), our proposed method can be efficiently implemented according to its definition; the corresponding computational complexity is at most $O(|{\mathcal{Y}}|mn)$ times the complexity of the selection rule.
In this section, we instantiate our general procedure beyond the small $|{\mathcal{Y}}|$ setting with concrete examples where special structures enable efficient computation. We focus on three classes of selection rules that can be of practical interest: selection using only the covariates, selection based on conformal p-values, and selection based on conformal prediction sets. When practitioners are willing to slightly modify the selection rule to improve computation efficiency, they may refer to Supplementary Section E.4 where we discuss extensions to simplify the construction of prediction sets by further splitting the calibration data.
@@ 4.1 Covariate-dependent selection rules
We first consider covariate-dependent selection rules, i.e., when ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is only a function of $\{X_{i}\}_{i=1}^{n+m}$. Under such rules, the reference set no longer depends on $y$; we shall suppress the dependence on $y$ and write ^$\widehat{{\mathcal{R}}}_{n+j}(y)\equiv\widehat{{\mathcal{R}}}_{n+j}$ throughout this subsection.
Here, ^$\widehat{{\mathcal{R}}}_{n+j}$ can be efficiently computed by looping over $i\in[n]$. The complete procedure for constructing $\widehat{C}_{\alpha,n+j}$ and $\widehat{C}^{\textnormal{rand}}_{\alpha,n+j}$ with an arbitrary covariate-dependent selection rule is summarized in Algorithm 1. Its overall computational complexity is $O(mn\cdot C_{\mathcal{S}})$, with $C_{\mathcal{S}}$ being the complexity of the selection process.
Algorithm 1 JOMI for arbitrary covariate-dependent selection rules
@@
[…]
Sample $U_{j}\sim\textnormal{Unif}[0,1]$.
[…]
By Theorem 1, the output of Algorithm 1 is valid as long as the selection rule does not rely on the ordering of the calibration points. This includes many commonly-used selection rules:
@@
    Top-K selection. The $K$ test units with the highest scores $S(X_{i})$ are selected, where ${\mathcal{S}}\colon{\mathcal{X}}\to\mathbb{R}$ is a pre-trained score function. For example, $S(X_{i})$ may be the predicted binding affinity for a drug with chemical structure $X_{i}$ or the predicted health risk of a patient with features $X_{i}$, and the drug discovery process or clinical system admits a fixed number of new units.
@@
    Selection based on joint quantiles. A unit $j$ is selected if its score $S(X_{n+j})$ surpasses the $q$-th quantile of both the calibration and test scores $\{S(X_{i})\}_{i=1}^{n+m}$. For instance, a scientist may be interested in toxicities of drug candidates in ${\mathcal{D}}_{\textnormal{test}}$ where those in ${\mathcal{D}}_{\textnormal{calib}}$ have been tested, but they only focus on drugs with highest predicted activities $S(X_{i})$ in the entire library.
@@
    Selection based on calibration quantiles. A test unit $j$ is selected if its score $S(X_{n+j})$ surpasses the $q$-th quantile of calibration scores $\{S(X_{i})\}_{i=1}^{n}$. This may happen when a doctor uses the predicted health risks of existing patients to determine a normal range, and picks test units anticipated to have a relatively extreme health risk.